# HyperSIGMA: Hyperspectral Intelligence Comprehension Foundation Model

### v1

@@ 4.1.3 Pre-training of Spectral SubNetwork
4.1.4 Implementation Details
Data Pre-processing Given the high-dimensional nature of HSIs, we perform dimensionality reduction by randomly selecting continuous channels to preserve the original spectral order and enhance the diversity of pre-training data. Since HSIs in the HyperGlobal-450K dataset are acquired by different sensors, our network cannot process images with varying numbers of channels. Therefore, the number of selected channels is consistent across all HSIs. Specifically, for any original HSI $\mathbf{X}_{HSI}\in\mathbb{R}^{H\times W\times L}=\{x_{1},x_{2},\ldots,x_{L}\}$ with $L$ channels, where $x_{i}\in\mathbb{R}^{H\times W}$ represents a channel, we randomly select continuous channels $\{x_{i},x_{i+1},\ldots,x_{i+C-1}\}$ by determining a value of $i$ that satisfies $1\leq i\leq L+1-C$. This results in $\mathbf{X}_{0}\in\mathbb{R}^{H\times W\times C}$, which is then processed by the network as described in Sec. 4.1.2 and 4.1.3. During pre-training, $C$ is fixed to ensure that different HSIs have the same number of channels. This channel selection strategy is illustrated in Fig. 5. In our implementation, $C$ is set to 100. Additionally, all HSIs are normalized by dividing by 4,000.
@@ 4.1.4 Implementation Details
Experimental Settings The mask ratio is a crucial hyperparameter in MIM algorithms, as it influences the difficulty of recovering masked regions and, consequently, the effectiveness of pre-training. A high mask ratio makes the reconstruction task overly challenging, potentially hindering restoration, while a low mask ratio may result in ineffective model weights due to the ease of the task. Thus, selecting an appropriate mask ratio $R$ is essential for ensuring pre-training quality. For spatial MAE, we adopted $R_{spat}=0.75$, a setting proven effective for both natural images [146] and remote sensing (RS) images [84]. In contrast, research on HSI channel masking and recovery is limited. For example, [103] independently determined relevant values for different datasets in HSI and LiDAR joint classification. Therefore, we conducted a series of ablation studies to identify a suitable $R_{spec}$ for spectral MAE.
Concretely, we pre-trained spectral subnetworks with different mask ratios ranging from 0.15 to 0.9, in intervals of 0.15, for 400 epochs with a batch size of 2,048. We used the AdamW optimizer [165], setting the learning rate to 0.00015 and weight decay to 0.05, with all other settings following those of MAE [146]. For convenience, the number of tokens $N_{spec}$ was set equal to $C$, mapping each channel to a token. After pre-training, we fine-tuned these networks on the Indian Pines (IP) and Pavia University (PU) datasets for HSI classification tasks. For IP, we used 10 samples per category for training, and for PU, 20 samples per category, with the remaining samples used for testing. Each input sample included all channels, and a 33$\times$33 patch centered on the target pixel was fed into the network. $N_{spec}$ remained set to 100. The network output, with a shape of $N_{spec}\times D$, was passed through an average pooling layer to reduce the dimensionality to $D$, followed by a linear layer for classification. We used cross-entropy loss for training and evaluated performance using overall accuracy (OA). The results, shown in Fig. 6, indicate that the best performance across all datasets was achieved with $R_{spec}=0.75$. Therefore, we used this mask ratio for the subsequent pre-training of spectral subnetworks.
TABLE I: Pre-training costs for various ViT versions in spatial and spectral subnetworks.
@@
After determining the mask ratios, we retrained the spatial and spectral subnetworks using the default settings as previously described. Following [146, 84], we extended the pre-training to 1,600 epochs for sufficient training. Given the image size of $\mathbf{X}_{0}$, the patch size $P$ for the spatial ViT is set to 8, resulting in $N_{spat}=\frac{H\cdot W}{P^{2}}=64$ tokens. The other training parameters of the spatial subnetwork, including batch size, learning rate, and weight decay, match those of the spectral subnetwork. We pre-trained different ViT versions - base, large, and huge - for both spatial and spectral networks, denoted as SpatViT and SpecViT, respectively. This allows HyperSIGMA’s model size to reach the billion level when using the ViT-Huge backbone for both subnetworks. Consequently, this pre-training required significant time and computational resources, as detailed in Table I. The differences between SpatViT and SpecViT within the same version mainly stem from the variations in embedding layers. Additionally, we employed Gradient Checkpointing when GPU memory was insufficient. All experiments were conducted on NVIDIA V100 GPUs.
4.2 Model Structure
4.2.1 Preliminaries of Vision Transformer
In this section, we briefly introduce ViTs [76] as they form the foundational structures of HyperSIGMA. Inspired by the transformer network used in NLP, where words are mapped into vectors, ViT operates by first splitting an image into non-overlapping patches. Each patch is then mapped to a 1-D token vector. The details of mapping spatial patches and spectral channels of HSI are detailed in Sec. 4.1.2, Sec. 4.1.3, and Sec. 4.1.4. These tokens, combined with positional embeddings $\mathbf{E}$, are processed by a series of transformer blocks $f=\left\{f_{1},f_{2},\cdots,f_{d}\right\}$, where $d$ is the network depth. This process can be formulated as:
$\mathbf{Z}$ $=LN(\mathbf{U}_{n}),\quad\mathbf{Z}\in\mathbb{R}^{N\times D}.$
$\begin{split}\mathbf{U}_{i-1}^{\prime}&=MHSA(LN(\mathbf{U}_{i-1}))+\mathbf{U}_{i-1},\\ \mathbf{U}_{i}&=FFN(LN(\mathbf{U}_{i-1}^{\prime}))+\mathbf{U}_{i-1}^{\prime}.\\ \end{split}$ (2)
$SA(\mathbf{U})=softmax(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{D^{\prime}}})\mathbf{V},$ (3)
$MHSA(\mathbf{U})=\left[Concat(SA_{1}(\mathbf{U}),\cdots,SA_{h}(\mathbf{U}))\mathbf{W}\right]^{T}.$ (4)
Here, $SA_{h}(\cdot)$ denotes the $h$-th SA head. $\mathbf{W}\in\mathbb{R}^{hD^{\prime}\times D}$ denotes the weight matrix of a linear layer to recover the original embedding dimension $D$. In practice, $D=h\cdot D^{\prime}$ for convenience.
4.2.2 Sparse Sampling Attention
In this section, we introduce the proposed SSA, designed to efficiently learn diverse contextual features by addressing the spatial and spectral redundancy of HSIs. Given $\mathbf{Q}=\{q_{1},\cdots,q_{N}\}$, $\mathbf{K}=\{k_{1},\cdots,k_{N}\}$, and $\mathbf{V}=\{v_{1},\cdots,v_{N}\}$, we predict $N_{p}$ offsets $[(\Delta x_{1},\Delta y_{1}),\cdots,(\Delta x_{N_{p}},\Delta y_{N_{p}})]$ using a linear layer $\bm{W}_{p}\in\mathbb{R}^{D^{\prime}\times 2N_{p}}$ for each query vector $q$ at coordinates $(c_{x},c_{y})$. Next, we sample new key $k^{\prime}$ and value $v^{\prime}$ from the original key and value matrices $\mathbf{K}$ and $\mathbf{V}$ at these positions using bilinear interpolation. Note that $\mathbf{K}$ and $\mathbf{V}$ have been reshaped into 2-D matrices of shape $\mathbb{R}^{H^{\prime}\times W^{\prime}\times D^{\prime}}$, where $N=H^{\prime}W^{\prime}$. This process can be formulated as:
$k^{\prime}_{j}=\sum_{(o_{x},o_{y})}$ $\max(0,1-|o_{x}-(c_{x}+\Delta x_{j})|)$ (5)
$\max(0,1-|o_{y}-(c_{y}+\Delta y_{j})|)\mathbf{K}[o_{x},o_{y},:],$
$v^{\prime}_{j}=\sum_{(o_{x},o_{y})}$ $\max(0,1-|o_{x}-(c_{x}+\Delta x_{j})|)$
$\max(0,1-|o_{y}-(c_{y}+\Delta y_{j})|)\mathbf{V}[o_{x},o_{y},:].$
$SSA(\mathbf{U})=softmax(\frac{\mathbf{Q}\otimes\mathbf{K}^{\prime}}{\sqrt{D^{\prime}}})\otimes\mathbf{V}^{\prime}.$ (6)
Figure 7: Comparison of different attention mechanisms: (a) Full SA [76], (b) WMHSA [70], (c) VSA [159], (d) RVSA [84], (e) DMHA [163], (f) SSA. Red and blue stars represent two queries, with corresponding colored lines (dots) indicating the regions of captured contexts (sampling positions). Green dots denote common areas shared by both queries. In DMHA, all queries share the same keys in the yellow region.
Comparison to Other Attention Methods Fig. 7 illustrates the differences between the proposed SSA and various attention methods. Full SA (Fig. 7 (a)) captures context across the entire input, while window-based attention methods [70, 159, 84] (Fig. 7 (b)-(d)) focus on rectangular regions. Deformable-DETR [162] introduces deformable attention, enhancing context diversity by using adaptive sampling points for each query, primarily for object detection. Inspired by this, DMHA [163] (Fig. 7 (e)) integrates deformable attention into vision transformer backbones, but all queries capture context from the same regions, even when different queries might need different contexts. Our SSA (Fig. 7 (f)) assigns a unique set of learnable sampling positions to each query, further enriching the context. Additionally, unlike deformable attention which uses linear layers to predict attention weights, SSA calculates attention weights through token interaction, enhancing the ability to capture inter-region relationships.
4.2.3 Structure of HyperSIGMA
Specifically, given a spatial feature $\mathbf{F}_{spat}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times D}$ and a spectral feature $\mathbf{F}_{spec}\in\mathbb{R}^{N_{spec}\times D}$, we first compress their dimensions using a linear layer to obtain $\mathbf{F}_{spat}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times D_{1}}$ and $\mathbf{F}_{spec}\in\mathbb{R}^{N_{spec}\times D_{1}}$. Next, we reduce the spatial dimension of $\mathbf{F}_{spec}$ through average aggregation to create a 1D vector $\mathbf{V}\in\mathbb{R}^{N_{spec}}$. Finally, we apply another linear layer to align the dimension of $\mathbf{V}$ with $\mathbf{F}_{spat}$. The spectral enhancement process is formulated as follows:
$\mathbf{F}^{\star}=(1+\mathbf{V})*\mathbf{F}_{spat},$ (7)
Since both $\mathbf{F}_{spat}$ and $\mathbf{F}^{\star}$ offer potential for interpreting HSIs, alongside HyperSIGMA, the spatial subnetwork termed SpatSIGMA also forms a foundational model for solving various high-level and low-level HSI tasks.
5 Experiments
In this section, we thoroughly evaluate HyperSIGMA’s performance on representative HSI high-level interpretation tasks: image classification (Sec. 5.1), target detection (Sec. 5.2), anomaly detection (Sec. 5.2), and change detection (Sec. 5.3), and low-level tasks: spectral unmixing (Sec. 5.4), image denoising (Sec. 5.5) and image super-resolution (Sec. 5.6). Additionally, we conduct experiments to examine HyperSIGMA’s scalability (Sec. 5.7.1), robustness (Sec. 5.7.2), cross-modal transferring capability (Sec. 5.7.3), and real-world applicability (Sec. 5.7.4).
Figure 9: Diagram illustrating the application of HyperSIGMA for hyperspectral image classification.
5.1 Hyperspectral Image Classification
5.1.1 Experiment Settings
Dataset Four widely used datasets are used, including Indian Pines [167], Pavia University [^3] https://www.ehu.eus/ccwintco/index.php/Hyperspectral_Remote_Sensing_Scenes, and HanChuan and HongHu scenes of the WHU-Hi dataset [168].
@@ 5.1.1 Experiment Settings
TABLE II: Fine-tuning accuracies of the spatial subnetwork with varying numbers of sampling points on the Indian Pines (IP) and Pavia University (PU) datasets. FA: Full Attention. NSP: Number of Sampling Points. Best and 2nd-best results are highlighted.
@@
TABLE III: Fine-tuning accuracies of the spatial subnetwork with various attentions on the Indian Pines (IP) and Pavia University (PU) datasets. FA: Full Attention; WS: Window Size; DR: Downsampling Rate; NSP: Number of Sampling Points. Best and 2nd-best results are highlighted.
Method WS DR NSP IP PU
FA - - - 58.25 77.54
WMHSA 4 - - 59.18 77.68
WMHSA 8 - - 56.79 78.84
VSA 4 - - 60.54 78.62
VSA 8 - - 60.72 80.37
RVSA 4 - - 62.74 79.49
RVSA 8 - - 60.61 79.74
DMHA - 4 - 47.06 69.93
DMHA - 8 - 45.78 67.98
SSA - - 8 63.92 82.00
TABLE IV: Fine-tuning accuracies of the spatial subnetwork using various pre-training weights on Indian Pines (IP) and Pavia University (PU) datasets. Best and 2nd-best results are highlighted.
@@
Implementation Details The classification task focuses on classifying all pixels in a hyperspectral image, similar to the semantic segmentation tasks in natural or aerial RGB images. This can be achieved through either patch-level or image-level classification [38]. Patch-level classification selects a patch centered around each pixel and classifies the center pixel based on this patch, while image-level classification corresponds directly to semantic segmentation. For computational efficiency, we primarily employ semantic segmentation unless otherwise specified. Fig. 9 illustrates the application of HyperSIGMA to this task. Spatial and spectral tokens are generated from the patch embedding layer. After processing through HyperSIGMA, four spatial feature maps $\bm{F}_{i}$ from the $i\cdot n/4$th layers ($i=1,\cdots,4$) are separately enhanced by the final spectral feature $\bm{V}$ via four distinct SEM modules. The output features from the SEM modules are then concatenated and processed by fully convolutional layers.
In the experiments, 33$\times$33 and 128$\times$128 patches are used for patch-level classification and segmentation, respectively, with a spatial patch size of 8. For the Indian Pines (IP), Pavia University (PU), HanChuan, and HongHu datasets, 10, 20, 50, and 50 labeled samples per class are randomly selected as the training set, while the remaining labeled samples form the test set. The model is trained for 500 epochs with a learning rate of 0.00006, and the number of tokens in the spectral branch is set to 100 ($N_{spec}=100$). Classification performance is assessed using overall accuracy (OA), average accuracy (AA), and the Kappa coefficient (Kappa). We compare HyperSIGMA against several classical and advanced classification and segmentation methods, including MSDN[169], SSFCN[128], FullyContNet[38], SpectralFormer[48], HSIC-FM[170], SSGRN [129], CSIL[171], and CLOLN[172].
5.1.2 Ablation Studies
To determine effective hyperparameter settings, we conduct a series of ablation experiments on the IP and PU datasets. We slightly modify the structure from Fig. 9. Only the features ($\bm{F}_{4}$) from the last layer are used directly for classification, omitting any subsequent layers. For simplicity, we use only the spatial subnetwork.
Figure 10: Predicted classification maps from various methods on the HongHu scene of the WHU-Hi dataset [168].
TABLE V: Performance (%) of various methods across four HSI classification datasets. Best and 2nd-best results are highlighted.
Metric MSDN [169] SSFCN[128] FullyContNet[38] SpectralFormer[48] HSIC-FM[170] SSGRN[129] CSILL[171] CLOLN[172] SpatSIGMA HyperSIGMA
@@ 5.1.2 Ablation Studies
OA 78.55 71.62 67.12 85.33 70.47 82.19 91.86 87.89 94.35 94.87
AA 73.06 66.51 73.81 83.20 71.03 87.31 94.64 78.25 88.10 87.18
Kappa 73.45 65.67 61.34 81.76 64.73 78.21 89.89 84.83 93.04 93.69
@@
Number of Sampling Points (NSP) NSP is a critical hyperparameter in SSA, significantly influencing contextual feature extraction. We performed experiments to analyze its impact, as shown in Table II. Here, “FA” denotes the use of full attention across all SpatSIGMA layers, while “SSA” indicates replacing some FAs with SSA (Sec. 4.2.3). SSA helps reduce data redundancy and generally enhances classification performance compared to FA, thereby validating the efficacy of our approach. Based on the results, we selected 8 sampling points for subsequent experiments.
Comparison of Different Attentions To further validate the superiority of SSA, we conducted a comparative analysis by replacing SSA with other attention mechanisms: VSA [159], RVSA [84], WMHSA [70], and DMHA [163]. VSA, RVSA, and WMHSA were evaluated with window sizes of 4 and 8, while DMHA used downsampling rates of 4 and 8. Table III summarizes the experimental results in terms of OA. SSA consistently achieved superior classification results. RVSA performed well, benefiting from its design tailored for RS overhead views. In contrast, DMHA exhibited poorer performance, primarily due to its neglect of hyperspectral objects beyond the global deformable region.
Comparison of Pre-training Weight Strategies Next, we compare various strategies for initializing model weights, including random initialization, pre-training on ImageNet [79, 80], MillionAID [138, 80], and HyperGlobal-450K. Results are summarized in Table IV. Loading weights pre-trained on different datasets occasionally led to performance degradation, likely due to domain gaps. Interestingly, pre-training with MillionAID, despite also being RS images, yielded lower accuracies, underscoring the uniqueness of hyperspectral data. Conversely, pre-training with HyperGlobal-450K significantly improved accuracy, highlighting its importance in HSI interpretation.
5.1.3 Results and Analyses
Table V presents the classification accuracies of different methods. Our models consistently outperform state-of-the-art approaches, even with limited labeled samples, achieving higher OA and Kappa scores. For example, on the HanChuan dataset, containing 16 visually similar agricultural categories, models were trained with 50 labeled samples per class (about 0.22% of the entire image). Despite these constraints, SpectralFormer achieves an OA of 82.60%, outperforming MSDN [169], HSIC-FM [170], FullyContNet [38], and SSFCN [128], albeit trailing behind our SpatSIGMA and HyperSIGMA. Notably, our approach also outperforms recent graph convolutional network [173] based method SSGRN [129], which achieves an OA of 90.43% but lags 4% behind HyperSIGMA. HyperSIGMA consistently outperforms SpatSIGMA across various scenes, leveraging spectral information for enhanced accuracy. However, our models show limitations in achieving high AA, particularly in classes with fewer pixels, suggesting challenges in capturing details of small objects due to larger patch sizes. Furthermore, qualitative evaluations through visualization of classification maps (Figure 10) on the HongHu dataset, demonstrate that our models effectively mitigate issues such as salt and pepper noise, over-smoothing, and misclassifications observed in comparative methods. These visualizations reveal clear and discriminative boundaries, highlighting the effectiveness of spectral enhancement in HyperSIGMA for refining boundaries. In summary, our models exhibit significant advancements in extracting robust and universal representations for HSI classification, demonstrating notable quantitative and qualitative performance improvements.
5.2 Hyperspectral Target and Anomaly Detection
@@
5.2.1 Experiment Settings
Dataset We adopt Mosaic [174] and AVIRIS [175] datasets for the HTD tasks, and Pavia [176] and Cri [177] for the HAD tasks.
Implementation Details Following HTD-ViT [178], we tackle the HTD task using coarse detection labels. Unlike HTD-ViT, which employs per-pixel classification, we draw inspiration from promptable segmentation [179] to design a segmentation network utilizing HyperSIGMA as a feature extractor. By reframing hyperspectral detection as a segmentation problem, we evaluate HyperSIGMA’s performance in both HTD and HAD tasks.
Specifically, we start by using CEM [180] for coarse detection [178]. Pixels with the highest 0.15% probabilities are labeled as pseudo detections, while the bottom 30% are set as the background. The remaining pixels are ignored during segmentation loss calculation. Fig. 11 illustrates the entire network framework. For segmentation, a linear layer matches the target spectrum $\mathbf{T}$’s dimension to $\mathbf{F}^{\star}$, i.e., from $\mathbb{R}^{C}$ ($C$ is the number of HSI channels) to $\mathbb{R}^{D_{1}}$. We then compute the similarities between each pixel of $\mathbf{F}^{\star}$ and $\mathbf{T}$:

### v2

@@ C.1 Implementation Details
Figure 14: Diagram illustrating the application of HyperSIGMA for HSI classification.
Figure 15: Predicted classification maps from various methods on the Indian Pines dataset [55].
Spectral Mask Ratio Concretely, we pre-trained spectral subnetworks with different mask ratios ranging from 0.15 to 0.9, in intervals of 0.15, for 400 epochs with a batch size of 2,048. We used the AdamW optimizer [218], setting the learning rate to 0.00015 and weight decay to 0.05, with all other settings following those of MAE [46]. For convenience, the number of tokens $N_{spec}$ was set equal to $C$, mapping each channel to a token. After pre-training, we fine-tuned these networks on the Indian Pines (IP) and Pavia University (PU) datasets for HSI classification tasks. For IP, we used 10 samples per category for training, and for PU, 20 samples per category, with the remaining samples used for testing. Each input sample included all channels, and a 33$\times$33 patch centered on the target pixel was fed into the network. $N_{spec}$ remained set to 100. The network output, with a shape of $N_{spec}\times D$, was passed through an average pooling layer to reduce the dimensionality to $D$, followed by a linear layer for classification. We used cross-entropy loss for training and evaluated performance using overall accuracy (OA). The results, shown in Fig. 13, indicate that the best performance across all datasets was achieved with $R_{spec}=0.75$. Therefore, we used this mask ratio for the subsequent pre-training of spectral subnetworks.
Pre-training Cost The pre-training required significant time and computational resources, as detailed in Table XIII. The differences between SpatViT and SpecViT within the same version mainly stem from the variations in embedding layers. Additionally, we employed Gradient Checkpointing when GPU memory was insufficient.
Appendix D Hyperspectral Image Classification
TABLE XIII: Pre-training costs for various ViT versions in spatial and spectral subnetworks.
@@ Appendix D Hyperspectral Image Classification
Figure 16: Predicted classification maps from various methods on the Pavia University dataset [^11] https://www.ehu.eus/ccwintco/index.php/Hyperspectral_Remote_Sensing_Scenes.
@@
TABLE XIV: Fine-tuning accuracies of the spatial subnetwork with varying numbers of sampling points on the Indian Pines (IP) and Pavia University (PU) datasets. FA: Full Attention. NSP: Number of Sampling Points. Best and 2nd-best results are highlighted.
@@
TABLE XV: Fine-tuning accuracies of the spatial subnetwork with various attentions on the Indian Pines (IP) and Pavia University (PU) datasets. FA: Full Attention; WS: Window Size; DR: Downsampling Rate; NAB: Number of Attention Buckets; NSP: Number of Sampling Points. Best and 2nd-best results are highlighted.
Method WS NAB DR NSP IP PU
FA [29] - - - - 58.25 77.54
WMHSA [28] 4 - - - 59.18 77.68
WMHSA [28] 8 - - - 56.79 78.84
VSA [159] 4 - - - 60.54 78.62
VSA [159] 8 - - - 60.72 80.37
RVSA [31] 4 - - - 62.74 79.49
RVSA [31] 8 - - - 60.61 79.74
NLSA [52] - 4 - - 58.98 79.09
NLSA [52] - 8 - - 61.08 80.54
DMHA [53] - - 4 - 47.06 69.93
DMHA [53] - - 8 - 45.78 67.98
SSA - - - 8 63.92 82.00
Figure 17: Predicted classification maps from various methods on the HanChuan dataset [56].
TABLE XVI: Fine-tuning accuracies of the spatial subnetwork using various pre-training weights on Indian Pines (IP) and Pavia University (PU) datasets. Best and 2nd-best results are highlighted.
@@
TABLE XVII: Performance (%) of various methods across different HSI classification datasets. Best and 2nd-best results are highlighted.
Metric MSDN [60] SSFCN[61] FullyContNet[59] SpectralFormer[20] HSIC-FM[62] SSGRN[63] CSIL[64] IDCN[65] CLOLN[66] SpatSIGMA HyperSIGMA
@@
OA 78.55 71.62 67.12 85.33 70.47 82.19 91.86 89.19 87.89 94.35 94.87
AA 73.06 66.51 73.81 83.20 71.03 87.31 94.64 88.15 78.25 88.10 87.18
Kappa 73.45 65.67 61.34 81.76 64.73 78.21 89.89 86.59 84.83 93.04 93.69
Houston
OA 72.18 72.39 51.07 77.21 54.43 68.62 66.11 85.34 85.95 87.33 86.80
AA 74.78 72.23 55.77 79.00 53.30 70.84 68.96 86.94 88.86 88.74 88.27
Kappa 69.93 70.12 47.00 75.32 50.80 65.99 63.21 84.11 84.78 86.30 85.72
ZY1-02D
OA 88.57 82.46 84.47 72.03 76.98 77.46 92.49 92.29 80.43 94.72 94.92
AA 88.13 80.61 83.69 63.06 74.75 82.28 91.31 92.89 73.88 95.70 93.95
Kappa 86.47 79.36 81.65 67.20 73.31 73.80 91.09 90.82 76.91 93.74 95.15
Figure 18: Predicted classification maps from various methods on the HongHu Wdataset [56].
D.1 Implementation Details
Figure 19: Predicted classification maps from various methods on the Houston dataset[57]
Figure 20: Predicted classification maps from various methods on the ZY1-O2D dataset [58].
Network Structure Fig. 14 illustrates the application of HyperSIGMA to this task. Spatial and spectral tokens are generated from the patch embedding layer. After processing through HyperSIGMA, four spatial feature maps $\bm{F}_{i}$ from the $i\cdot n/4$th layers ($i=1,\cdots,4$) are separately upsampled to the input size and enhanced by the final spectral feature $\bm{V}$ via four distinct SEM modules. The output features from the SEM modules are then concatenated and processed by fully convolutional layers.
Experimental Details In the experiments, 33$\times$33 patches are used for patch-level classification. For patch-level segmentation, besides 32$\times$32 patches and a spatial patch size of 2 are adopted for the Pavia University dataset with the consideration of tiny objects, other datasets apply 128$\times$128 patches, where the spatial patch size is set to 2 for Houston and ZY datasets, and 8 for other datasets. For the Indian Pines, Pavia University, HanChuan, HongHu, and ZY1-02D datasets, 10, 20, 50, 50, and 20 labeled samples per class are randomly selected as the training set, the remaining labeled samples form the test set, while the Houston dataset follows the official train/test split. The model is trained for 500 epochs with a learning rate of 0.00006, and the number of tokens in the spectral branch is set to 100 ($N_{spec}=100$).
D.2 Ablation Studies
Network Structure We slightly modify the structure from Fig. 14. Only the features ($\bm{F}_{4}$) from the last layer are used directly for classification, omitting any subsequent layers.
@@ D.2 Ablation Studies
Number of Sampling Points (NSP) NSP is a critical hyperparameter in SSA, significantly influencing contextual feature extraction. We performed experiments to analyze its impact, as shown in Table XIV. Here, “FA” denotes the use of full attention across all SpatSIGMA layers, while “SSA” indicates replacing some FAs with SSA. SSA helps reduce data redundancy and generally enhances classification performance compared to FA, thereby validating the efficacy of our approach. Based on the results, we selected 8 sampling points for subsequent experiments.
@@
Figure 22: Diagram illustrating the application of HyperSIGMA for hyperspectral anomaly detection.
Figure 23: Detection probability maps predicted by various methods for the HTD tasks. Top row: Mosaic [68]. Middle row: AVIRIS[69]. Bottom row: Renourban [70].
Comparison of Different Attentions To further validate the superiority of SSA, we conducted a comparative analysis by replacing SSA with other attention mechanisms: VSA [159], RVSA [31], NLSA[52], WMHSA [28], and DMHA [53]. VSA, RVSA, and WMHSA were evaluated with window sizes of 4 and 8, NLSA separately adopted 4 and 8 attention buckets, while DMHA used downsampling rates of 4 and 8. Table XV summarizes the experimental results in terms of OA. SSA consistently achieved superior classification results. RVSA performed well, benefiting from its design tailored for RS overhead views. NLSA also obtained competitive accuracies by capturing locality-sensitive contexts. In contrast, DMHA underperformed due to its limited focus on hyperspectral objects within the global deformable region.
Comparison of Pre-training Weight Strategies Next, we compare various strategies for initializing model weights, including random initialization, pre-training on ImageNet [160, 37], MillionAID [162, 37], and HyperGlobal-450K. Results are summarized in Table XVI. Loading weights pre-trained on different datasets occasionally led to performance degradation, likely due to domain gaps. Interestingly, pre-training with MillionAID, despite also being RS images, yielded lower accuracies, underscoring the uniqueness of hyperspectral data. Conversely, pre-training with HyperGlobal-450K significantly improved accuracy, highlighting its importance in HSI interpretation.
D.3 More Results
Additional Metrics Besides OA, we provide more metrics of the classification result, including average accuracy (AA) and the Kappa coefficient (Kappa), as shown in Table XVII. It can be observed that our models show limitations in achieving high AA on the first five datasets, particularly in classes with fewer pixels, suggesting challenges in capturing details of small objects due to large patch sizes.
Visualization Furthermore, qualitative evaluations through visualization of classification maps (Fig. 15-20) for the Indian Pines, Pavia University, HanChuan, HongHu, Houston and ZY1-02D datasets, demonstrate that our models effectively mitigate issues such as salt and pepper noise, over-smoothing, and misclassifications observed in comparative methods. These visualizations reveal clear and discriminative boundaries, highlighting the effectiveness of spectral enhancement in HyperSIGMA for refining boundaries.
Appendix E Hyperspectral Target and Anomaly Detection
Figure 24: Detection probability maps predicted by various methods for the HAD tasks. Top row: Pavia[71]. Middle row: Cri[72]. Bottom row: Viareggio [73].
Figure 25: ROC curves for various methods applied to the HTD and HAD tasks.
E.1 Implementation Details
Network Framework Specifically, we start by using CEM [76] for coarse detection [74]. Pixels with the highest 1.5% probabilities are labeled as pseudo detections, while the bottom 30% are set as the background by following [74]. The remaining pixels are ignored during segmentation loss calculation. Fig. 22 illustrates the entire network framework. For segmentation, a linear layer matches the target spectrum $\mathbf{T}$’s dimension to $\mathbf{F}^{\star}$, i.e., from $\mathbb{R}^{C}$ ($C$ is the number of HSI channels) to $\mathbb{R}^{D_{1}}$. We then compute the similarities between each pixel of $\mathbf{F}^{\star}$ and $\mathbf{T}$: